OpenStudy (anonymous):

How many solutions does this system of equations have? 2x - 2y = 8 2x + 3y = 3

OpenStudy (campbell_st):

one ordered pair

OpenStudy (anonymous):

How many solutions does this system of equations have? 2x - 2y = 8=y=x-4 2(x-y)=8 2x-2y-8=0

OpenStudy (anonymous):

y=x-4

OpenStudy (anonymous):

it only has 1 correct?

OpenStudy (anonymous):

It only has 1 solution. You can tell because if you get both equations into slope-intercept form (that is, solve for y in both equations), they each have a different slope. With different slopes, you've only got 1 solution.

OpenStudy (anonymous):

2x + 3y = 3=y=12x/3 2x+3y-3=0 x=3n y=1-2n y=1/3(3-2x)

OpenStudy (campbell_st):

the solution is the order pair or point of intersection of the 2 lines...